### v3

# Quantization Error Propagation: Revisiting Layer-Wise Post-Training Quantization

Abstract
Layer-wise PTQ is a promising technique for compressing large language models (LLMs), due to its simplicity and effectiveness without requiring retraining. However, recent progress in this area is saturating, underscoring the need to revisit its core limitations and explore further improvements. We address this challenge by identifying a key limitation of existing layer-wise PTQ methods: the growth of quantization errors across layers significantly degrades performance, particularly in low-bit regimes. To address this fundamental issue, we propose Quantization Error Propagation (QEP), a general, lightweight, and scalable framework that enhances layer-wise PTQ by explicitly propagating quantization errors and compensating for accumulated errors. QEP also offers a tunable propagation mechanism that prevents overfitting and controls computational overhead, enabling the framework to adapt to various architectures and resource budgets. Extensive experiments on several LLMs demonstrate that QEP-enhanced layer-wise PTQ achieves substantially higher accuracy than existing methods. Notably, the gains are most pronounced in the extremely low-bit quantization regime.
1 Introduction
Large Language Models (LLMs) have achieved impressive performance in various natural language processing tasks, including open-ended text generation, multi-step reasoning, and dialogue modeling. Notable examples include ChatGPT (achiam2023gpt) and the Llama family (touvron2023llama; grattafiori2024llama). However, deploying LLMs cost-effectively remains difficult because of their substantial memory usage and computational demands (chen2023frugalgpt). This limitation is especially critical for edge computing and latency-sensitive applications. To address these challenges, a wide range of model compression techniques, such as quantization (lang2024comprehensive; gong2024survey), pruning (wang2024model; cheng2024survey), low-rank approximation (yang2024low; hu2022lora), and knowledge distillation (xu2024survey; yang2024survey), have been explored.
Among these methods, layer-wise post-training quantization (PTQ) has emerged as a practical and widely used approach for large-scale LLMs (frantar2022gptq; lin2024awq; yao2022zeroquant; chee2023quip). Unlike block-wise PTQ (tseng2024quip; shao2023omniquant), global fine-tuning (egiazarian2024extreme; tseng2024quip), quantization-aware training (QAT) (xu2024onebit; wang2023bitnet; liu2023llm), and all of which require heavy retraining and backpropagation, layer-wise PTQ quantizes model parameters layer-by-layer without retraining or backpropagation, resulting in significantly lower computational and memory demands. Despite its simplicity, layer-wise PTQ effectively preserves model quality even at lower bit widths (frantar2022gptq; lin2024awq; chee2023quip). As a result, layer-wise PTQ is increasingly adopted in real-world applications due to its efficient quantization, reduced computational cost, and broader compatibility with large-scale LLMs, varying bit widths, and diverse quantization strategies.
Despite significant progress in layer-wise PTQ, advancements in this area are saturating (malinovskii2024pv). This study aims to push the performance boundaries of layer-wise PTQ by revisiting its core design strategy. This study begins by identifying a fundamental limitation of existing layer-wise PTQ approaches. These approaches do not adequately account for the propagation of quantization errors across layers. Quantization errors accumulate significantly, leading to a degradation in overall model performance, especially in low-bit settings. This represents a key bottleneck for the practical deployment of layer-wise PTQ in large-scale LLMs.
Figure 1: WikiText‑2 perplexity comparison across Llama‑2 models (7B-70B) quantized to INT‑4, INT‑3, and INT‑2, employing RTN, GPTQ, AWQ, and QuIP methods. Solid bars indicate PTQ with QEP; border bars represent PTQ without QEP. Truncated bars indicate perplexities exceeding axis limits. QEP consistently reduces perplexity, with greater improvements observed at lower bitwidths and smaller model sizes. See Section 6 for detailed settings and results.
To address this issue, we propose Quantization Error Propagation (QEP), a general and computationally efficient framework that enhances the performance of layer-wise PTQ methods. QEP modifies the layer-wise optimization objective to propagate and compensate for accumulated quantization errors, while maintaining computational complexity comparable to existing layer-wise PTQ methods. Furthermore, we introduce a tunable propagation mechanism whose adjustable propagation strength prevents overfitting, a known issue previously observed in GPTQ (lin2024awq). This mechanism also enables adaptive control over computational overhead, especially in parameter-heavy components such as MLP blocks. Notably, the enhancement of QEP is orthogonal to existing PTQ methods and can be seamlessly integrated with any layer-wise PTQ pipeline.
Extensive experiments on several LLMs across various bit-width settings show that QEP significantly enhances layer-wise PTQ methods, including GPTQ (frantar2022gptq), AWQ (lin2024awq), QuIP (chee2023quip), as shown in Figure 1. These improvements are particularly pronounced in extreme low-bit regimes, such as 2-bit quantization, where standard layer-wise PTQ methods typically degrade significantly.
2 Related Work
Quantization techniques primarily include data-free PTQ (dettmers2023case), layer-wise PTQ (frantar2022gptq; lin2024awq; chee2023quip), block-wise PTQ (tseng2024quip; shao2023omniquant), global fine-tuning PTQ (egiazarian2024extreme; tseng2024quip), and QAT (xu2024onebit; wang2023bitnet; liu2023llm). Among these methods, weight-only layer-wise PTQ has become especially popular for large-scale models because of its computational efficiency and strong performance (frantar2022gptq; lin2024awq; chee2023quip). Recent benchmarking further highlights that most PTQ advances specifically target layer-wise methods (zhao2025benchmarking). Following the taxonomy in (zhao2025benchmarking), we outline three distinct approaches and recent developments.
Compensation-based layer-wise PTQ
This category, pioneered by GPTQ (frantar2022gptq), uses a sequential quantization strategy, in which model weights are quantized based on the Hessian computed from a calibration dataset, while compensating for subsequent unquantized weights. Several studies refined the compensation mechanism by improving update rules (behdin2023quantease), integrating nonlinear quantization schemes (liu2024vptq), employing adaptive grid selection (zhang2024leanquant), and using block-wise optimization (guan2024aptq).
Rotation-based layer-wise PTQ
A second promising direction, advanced by QuIP (chee2023quip), involves preprocessing weights through structured rotation matrices to more uniformly redistribute weight magnitudes. This approach was improved by randomized Hadamard transforms and block-wise and global fine-tuning optimization (tseng2024quip). Learning-based methods to determine rotation matrices have also been introduced (liu2024spinquant). This rotation-based strategy has also been extended to activation quantization (ashkboos2024quarot).
Salience-based layer-wise PTQ
Other approaches focus on identifying and preserving salient weights, often using mixed-precision quantization frameworks (dettmers2022gpt3; dettmers2023spqr; shang2023pb). Although mixed-precision methods usually add complexity due to various data types, AWQ (lin2024awq) mitigates these implementation difficulties. AWQ strategically employs a global scaling mechanism to align salient weights with the quantization grid better, simplifying deployment while maintaining high accuracy.
Recent advances in layer-wise PTQ have mainly focused on nonlinear quantization and block-wise and global fine-tuning extensions; however, the fundamental layer-wise optimization has remained largely unchanged since GPTQ (frantar2022gptq). This study revisits this foundational strategy, identifies its key limitations, and proposes improvements, demonstrating performance gains on the fundamental benchmarks such as GPTQ (frantar2022gptq), QuIP (chee2023quip), and AWQ (lin2024awq). Therefore, our contributions complement and are orthogonal to recent advancements, such as nonlinear quantization and structured extensions.
3 Background
Post-training quantization
Post-training quantization (PTQ) is a technique that converts the parameters of pre-trained models into discrete quantized representations. Formally, let ${\bm{W}}_{l}\in{\mathbb{R}}^{n_{l}\times d_{l}}$ denote the pre-trained weight matrix associated with the $l$-th linear operation. Note that the index $l$ specifically refers to individual linear transformations rather than entire transformer blocks. The objective of PTQ is to find a quantized approximation $\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}$ that closely approximates the behavior of the original model, preserving performance while reducing computational costs and memory usage. The set $\mathbb{Q}$$\subset{\mathbb{R}}$ denotes the discrete quantization domain, which is represented as a finite set of $2^{b}$ distinct quantization levels, referred to as a $b$-bit quantization scheme. To achieve accurate quantization, many approaches leverage a small calibration dataset. Specifically, given a calibration dataset ${\bm{X}}\in{\mathbb{R}}^{d_{1}\times m}$ consisting of $m$ samples, these methods aim to find optimal quantized parameters $\widehat{{\bm{W}}}_{l}$ that minimizes the deviation from the performance of the original model.
Layer-wise PTQ
Layer-wise PTQ has emerged as a promising framework (frantar2022gptq; frantar2022optimal) for compressing large-scale LLMs. Recent advancements in this area have significantly reduced the computational overhead and memory requirements of deploying LLMs. Despite methodological differences, existing layer-wise PTQ approaches typically follow a shared sequential quantization scheme, processing each layer independently and sequentially from the input layer toward the output layer.
Formally, these techniques quantize the model parameters $\{{\bm{W}}_{l}\}_{l=1}^{L}$ by solving the following layer-wise independent optimization problem:
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}}\left\|{\bm{W}}_{l}\mathsf{X}_{l}-\widehat{{\bm{W}}}_{l}\mathsf{X}_{l}\right\|_{F}^{2},$ (1)
where $\mathsf{X}_{l}$ denotes the input activations to the $l$-th layer. This quantization proceeds sequentially from $l=1$ toward the output layers. Due to the quadratic form of the reconstruction objective, the associated Hessian, $\mathsf{H}_{l}\coloneqq\mathsf{X}_{l}\mathsf{X}_{l}^{\top}$, can be efficiently precomputed and cached for reuse in subsequent optimization steps, improving computational efficiency in practice.
Existing PTQ methods typically use one of two possible forms for activation inputs $\mathsf{X}_{l}$: Either quantized activations ${\bm{X}}_{l}$, obtained by forward propagating the calibration dataset through previously quantized weights $\{\widehat{{\bm{W}}}_{1},\ldots,\widehat{{\bm{W}}}_{l-1}\}$, or full-precision activations ${\bm{X}}_{l}$, resulting from forward propagation through the original, unquantized weights $\{{\bm{W}}_{1},\ldots,{\bm{W}}_{l-1}\}$. There is no consensus among existing PTQ methods (frantar2022gptq; lin2024awq; chee2023quip) regarding whether quantized or full-precision activations produce better quantization outcomes.
Leading layer-wise PTQ methods use distinct optimization strategies to approximate the behavior of the original model while adhering to the foundational sequential layer-wise framework in Eq. (1). GPTQ (frantar2022gptq), for example, uses quantized activations, $\mathsf{X}_{l}=\widehat{{\bm{X}}}_{l}$, and quantizes parameters row-wise by sequentially minimizing reconstruction error and correcting residuals in the remaining unquantized entries until each row is fully quantized. AWQ (lin2024awq) uses original activations, $\mathsf{X}_{l}={\bm{X}}_{l}$, and identifies a small subset of salient weights whose magnitudes significantly influence the layer outputs, subsequently rescaling these weights before quantization.
4 Bottleneck: Quantization Error Accumulation and Growth
Figure 2: Accumulation and growth of quantization errors across layers in a partially quantized Llama2-7B model (touvron2023llama). The first $10$ Transformer blocks are quantized using standard RTN (BASE) and QEP-enhanced RTN (With QEP), while the remaining Transformer blocks after the 1010th remain at full precision. The plot shows the squared Frobenius norm $\Delta_{m}$, defined in Eq. (2), between the original and partially quantized outputs at each Transformer block $m$.
To motivate our proposed approach, this section first revisits the core layer-wise optimization formulation given by Eq. (1), emphasizing its key limitation: The accumulation and growth of quantization errors across layers significantly degrade the performance. We investigate this phenomenon using experiments conducted on the pre-trained Llama-2-7B model (touvron2023llama). Specifically, we quantize only the first $10$ Transformer blocks (vaswani2017attention), while keeping all subsequent blocks in full precision. To quantify the propagation and accumulation of the errors, we measure the discrepancy between fully precise and partially quantized outputs at each block using a calibration dataset. Let $\mathtt{TransBlock}_{m}(\cdot)$ denote the original full-precision $m$-th Transformer block, and $\widehat{\mathtt{TransBlock}}_{m}(\cdot)$ denote its quantized counterpart. We evaluate the following metric at the $m$-th block:
$\Delta_{m}=\left\|f_{m}({\bm{X}})-\widehat{f}_{m}({\bm{X}})\right\|_{F}^{2},$ (2)
$f_{m}({\bm{X}})\coloneqq\mathtt{TransBlock}_{m}\circ\cdots\circ\mathtt{TransBlock}_{n+1}\circ\mathtt{TransBlock}_{n}\circ\cdots\circ\mathtt{TransBlock}_{1}({\bm{X}}),$ (3)
$\widehat{f}_{m}({\bm{X}})\coloneqq\mathtt{TransBlock}_{m}\circ\cdots\circ\mathtt{TransBlock}_{n+1}\circ\widehat{\mathtt{TransBlock}}_{n}\circ\cdots\circ\widehat{\mathtt{TransBlock}}_{1}({\bm{X}}).$ (4)
This experiment sets $n=10$. Figure 2 shows an approximately exponential accumulation or errors within the quantized layer, as well as an error growth that persists in the unquantized layers. This growth occurs due to the layer-wise independent quantization approach described in Eq. (1), which neither accounts for quantization error propagated from previous layers nor corrects previously accumulated errors, thus exacerbating error growth in subsequent unquantized layers. The exponential accumulation of quantization errors observed empirically can also be theoretically explained under mild conditions, as detailed in Appendix B.2. Therefore, instead of treating layer-wise quantization as a series of independent optimization problems, it is essential to reformulate the original layer-wise optimization presented in Eq. (1) to mitigate error accumulation and growth.
5 QEP: Quantization Error Propagation
Existing layer-wise independent PTQ has inherent limitations, particularly the accumulation and growth of quantization errors discussed in Section 4. To address these limitations, we introduce Quantization Error Propagation (QEP), a general, lightweight, and scalable framework that improves layer-wise PTQ by propagating quantization errors. In subsequent sections, we provide theoretical evidence showing that QEP effectively reduces quantization errors.
5.1 Problem Reformulation
We reformulate the layer-wise independent optimization strategy presented in Eq. (1) to propagate quantization errors across layers effectively. Instead of minimizing output differences based on shared input activations $\mathsf{X}_{l}$, our reformulation directly minimizes the discrepancy between full-precision and quantized outputs, each computed using their respective upstream inputs. Formally, for each layer $l$, we optimize the discrete quantized weight matrix $\widehat{{\bm{W}}}_{l}$ as follows:
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}}\left\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\right\|_{F}^{2}.$ (5)
This objective ensures that the quantized weights $\widehat{{\bm{W}}}_{l}$ are optimized not only to independently approximate the full-precision weights ${\bm{W}}_{l}$ but also to counteract and compensate for the cumulative quantization errors introduced by previous layers. In contrast to the existing objective in Eq. (1), where the trivial optimal solution is $\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}$ if ${\bm{W}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}$, the optimal solution under the formulation in Eq. (5) is generally $\widehat{{\bm{W}}}_{l}\neq{\bm{W}}_{l}$, explicitly enabling error correction and accounting for accumulated quantization errors.
Although the modification from Eq. (1) seems straightforward, Eq. (5) inherently breaks the key structural simplification that facilitates efficient quantization in existing PTQ frameworks. Specifically, the optimization in Eq. (5) no longer solely depends on the Hessian matrix $\mathsf{H}_{l}$, thereby preventing the direct use of existing Hessian-based acceleration methods for quantization. In the following section, we address this challenge by proposing a practical and efficient weight correction scheme that overcomes this limitation while retaining the advantages of our error-propagation approach.
5.2 Weight Correction
To efficiently perform quantization by the objective in Eq. (5) as in existing layer-wise PTQ methods, we relax the discrete feasible set to a continuous domain, leading to the following proposition.
Proposition 5.1.
Assume that the matrix $\widehat{{\bm{H}}}_{l}$ is invertible. Then, after relaxing the discrete feasible set ${\mathbb{Q}}^{n_{l}\times d_{l}}$ into the continuous domain ${\mathbb{R}}^{n_{l}\times d_{l}}$, the optimal solution ${\bm{W}}_{l}^{\ast}$ is given by the following closed-form expression:
${\bm{W}}_{l}^{\ast}\coloneqq{\bm{W}}_{l}+{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1}=\operatorname*{argmin}_{\widehat{{\bm{W}}}_{l}\in{\mathbb{R}}^{n_{l}\times d_{l}}}\left\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\right\|_{F}^{2},$ (6)
where ${\bm{\delta}}_{l}\coloneqq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$ represents the accumulated quantization error from proceeding layers, $\widehat{{\bm{H}}}_{l}\coloneqq\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$ denotes the empirical Hessian constructed from quantized activations.
The proof of Proposition 5.1 is provided in Appendix B.1. Proposition 5.1 highlights an important distinction from the existing formulation given by Eq. (1). Specifically, when upstream quantization introduces non-negligible errors, i.e., ${\bm{\delta}}_{l-1}\neq 0$, the optimal quantized weights differ from straightforward approximations of the original weights ${\bm{W}}_{l}$. Instead, the optimal solution explicitly includes a correction term that compensating for accumulated quantization errors.
This corrected weight enables us to reformulate the equivalent optimization objective within the original discrete set $\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}$ as follows:
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}}\left\|{\bm{W}}_{l}^{\ast}\widehat{{\bm{X}}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\right\|_{F}^{2}.$ (7)
This objective shares the same structure as Eq. (1), with ${\bm{W}}_{l}$ replaced by its corrected counterpart ${\bm{W}}_{l}^{\ast}$. This reformulation restores the quadratic structure found in Eq. (1), facilitating efficient optimization through the Hessian matrix $\mathsf{H}_{l}=\widehat{{\bm{H}}}_{l}$. The structure of Eq. (7) allows for seamless integration with various existing layer-wise PTQ methods, as discussed in Section2. Furthermore, the proposed layer-wise quantization formulation in Eq. (5) formally guarantees improved quantization accuracy compared to the existing layer-wise independent PTQ defined in Eq. (1). Specifically, we establish the following theoretical result:
Theorem 5.2 (Informal).
Consider an $L$-layer neural network defined by:
$f_{{\bm{\theta}}}(X)=\sigma_{L}({\bm{W}}_{L}\sigma_{L-1}({\bm{W}}_{L-1}\cdots\sigma_{2}({\bm{W}}_{2}\sigma_{1}({\bm{W}}_{1}{\bm{X}}))\cdots)),$ (8)
where each activation function $\sigma_{l}$ is Lipschitz continuous and $\bm{\theta}$} denotes the set of all full-precision parameters $\{{\bm{W}}_{l}\}_{l=1}^{L}$. The output quantization error of the proposed quantization method defined in Eq. (5) is bounded by that of the existing layer-wise PTQ defined in Eq. (1):
$\left\|f_{\theta}({\bm{X}})-f_{\widehat{{\bm{\theta}}}_{\mathrm{QEP}}}({\bm{X}})\right\|_{F}\leq\left\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{{\bm{\theta}}}_{\mathrm{BASE}}}({\bm{X}})\right\|_{F}.$ (9)
where $\widehat{{\bm{\theta}}}_{\mathrm{QEP}}$ and $\widehat{{\bm{\theta}}}_{\mathrm{BASE}}$ denote the sets of parameters quantized by the objective in Eq. (5) and the base PTQ method by the objective in Eq. (1), respectively.
Explicit conditions and detailed proof are provided in Appendix B.3. The additional computational overhead arises solely from computing the correction term ${\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$, since computing the Hessian inverse $\widehat{{\bm{H}}}_{l}^{-1}$ remains unchanged from existing layer-wise independent PTQ. As empirically demonstrated in Section 6.1, this additional computation requires significantly less runtime compared to the quantization processes of layer-wise PTQ methods, even for large-scale LLMs, due to the tunable mechanism described in the next section.
5.3 Controlling Propagation Strength
Although solving Eq. (7) effectively reduces the accumulation of quantization error, it can lead to overfitting. This issue is particularly pronounced when the calibration dataset is small and insufficiently representative of the target task, or when the model includes blocks with a large number of parameters such as the MLP blocks commonly found in transformer architectures, causing the correction to overfit the calibration dataset.
To address this issue, we introduce a tunable propagation mechanism that generalizes the correction term using a scaling parameter $\alpha_{l}\in[0,1]$:
${\bm{W}}_{l}^{\ast}(\alpha_{l})={\bm{W}}_{l}+\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1}.$ (10)
Here, setting $\alpha_{l}=1$ recovers original fully-corrected case presented in Eq. (6), whereas setting $\alpha_{l}=0$ corresponds to the existing approach in Eq. (1) under the setting that $\mathsf{X}_{l}=\widehat{{\bm{X}}}_{l}$. This tunable correction mechanism relates to the following regularization optimization:
Proposition 5.3.
The parameter $\alpha_{l}$ corresponds to the regularization parameter $\lambda$ in the following optimization problem:
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}^{2}+\lambda_{l}\|{\bm{W}}_{l}-\widehat{{\bm{W}}}_{l}\|_{F}^{2},~~\lambda_{l}\in{\mathbb{R}}_{+}.$ (11)
Specifically, as $\alpha_{l}$ increases from 0 to $1$, the corresponding parameter $\lambda_{l}$ decreases from $\infty$ to 0.
The derivation is provided in Appendix B.4. Additionally, the following proposition is established.
Proposition 5.4.
Under the same assumptions in Theorem 5.2, the output quantization error of the method employing QEP with parameter $\{\alpha_{l}\}_{l=1}^{L}$ decreases monotonically as each $\alpha_{l}$ approaches $1$.
Explicit conditions and comprehensive proofs of this proposition are provided in Appendix B.3. Consequently, the parameter $\alpha_{l}$ effectively controls overfitting, analogous to regularization techniques, and importantly provides a systematic way to balance overfitting and underfitting in layer-wise PTQ methods. Indeed, this parameter is crucial for preventing overfitting, especially in MLP blocks, which contain more parameters than other blocks.
Furthermore, in large-scale LLMs, the high-dimensional activations in MLP layers often result in computationally expensive correction terms. In these cases, selectively setting $\alpha_{l}=0$ for specific layers eliminates the computational cost of the correction term and acts as implicit regularization, potentially improving generalization. Therefore, appropriately setting $\alpha_{l}=0$ can reduce the correction time by approximately one-third and one-half. Developing adaptive strategies for layer-wise, data-aware, or resource-efficient tuning of $\alpha_{l}$ is a promising direction for future research. In the following, we refer to the overall approach, including the tunable mechanism controlled by $\{\alpha_{l}\}_{l=1}^{L}$, as Quantization Error Propagation (QEP).
6 Experiments
We conduct experiments to validate the effectiveness of QEP in improving the performance of layer-wise PTQ relative to existing methods.
Baselines
We use representative layer-wise PTQ methods based on linear quantization such as round-to-nearest (RTN) (frantar2022gptq; dettmers2023case), GPTQ (frantar2022gptq), AWQ (lin2024awq), and QuIP (chee2023quip). Although previous studies have explored extensions, such as non-linear and block-wise quantization, as discussed in Section 2, these techniques are orthogonal to the core improvement introduced by QEP. Therefore, to isolate and emphasize the impact of QEP, we focus on these representative layer-wise PTQ methods.
Quantization
This study focuses on weight-only quantization schemes, specifically per-channel and group-wise quantization, which have recently shown superior trade-offs between efficiency and accuracy (dettmers2023case; frantar2022gptq; lin2024awq). The main text evaluates per-channel quantization under INT4, INT3, and INT2 precision settings. Due to space constraints, detailed results for group-wise quantization are presented in Appendix D. For the propagation strength parameter $\alpha_{l}$, we adopt a representative default value of $\alpha_{l}=\nicefrac{{1}}{{2}}$ for all layers, except for the MLP layers in the Llama-2 70B model, for which we set $\alpha_{l}=0$. Tuning $\alpha_{l}$ can further improve performance but is beyond the scope of this study and is left for future work.
Datasets
Following previous studies, we evaluate the Hessian matrix using the same default calibration datasets used in their original implementations. Specifically, GPTQ and QuIP use the C4 dataset (frantar2022gptq) for calibration, while AWQ uses the Pile dataset (gao2020pile). Following frantar2022gptq, we evaluate the correction term in Eq. (6) using $128$ randomly sampled segments of $2048$ tokens each from the C4 dataset(raffel2020exploring), which consists of web-crawled text excerpts.
Models
Following lin2024awq; frantar2022gptq, we evaluate our method on recent popular LLMs, namely the Llama-2 and Llama-3 model families (touvron2023llama), with size ranging from 7B to 70B parameters, as well as Mistral-7 B (jiang2024identifying). These models demonstrate superior performance compared to other open-source LLMs (zhang2022opt; scao2022bloom) and have become widely adopted as foundational models for numerous derivative open-source models (taori2023stanfordalpaca; vicuna2023).
Evaluations
Following established evaluation protocols from prior studies (dettmers2022gpt3; xiao2023smoothquant; frantar2022gptq; dettmers2023case; yao2022zeroquant), we evaluate the quantized LLMs using the perplexity (PPL) on WikiText2 (merity2016pointer), Penn Treebank (PTB) (marcus-etal-1994-penn), and C4 (raffel2020exploring), and zero-shot accuracy on benchmarks including ARC Easy (ArcE) (Boratko2018), PiQA (bisk2020piqa), and StoryCloze (SC) (mostafazadeh-etal-2016-corpus). Due to space limitations, detailed results for each dataset are provided in Appendix D. All experiments are conducted using a single NVIDIA V100 GPU.
6.1 Results
Perplexity
Table 1: Evaluation of perplexities ($\downarrow$) for Llama models on WikiText-2 under various layer-wise PTQ methods and bitwidths.
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B Llama-3-8B Mistral-7B
FP16 - - 5.472 4.883 3.319 6.137 5.255
INT4 RTN 6.116 5.206 3.672 8.540 5.997
6.017 5.165 3.621 8.021 5.877
GPTQ 6.083 5.167 3.594 147.912 5.643
5.933 5.127 3.576 9.509 5.528
AWQ 5.831 5.064 3.484 7.108 5.716
5.756 5.041 3.479 6.981 5.636
QuIP 8.434 5.137 3.826 6.998 11.109
5.753 5.034 3.485 6.650 5.479
INT3 RTN 539.866 10.688 7.530 2276.227 29.390
17.309 7.458 5.648 86.430 10.241
GPTQ 10.881 6.632 4.860 64.457 8.247
7.898 6.245 4.102 18.845 7.347
AWQ 15.299 6.448 4.362 11.802 7.902
11.131 6.092 4.103 10.713 7.169
QuIP 12.048 5.503 4.135 8.288 7.108
6.154 5.347 3.813 7.703 5.842
INT2 RTN 17783.918 51152.832 26077.172 1437176.750 78488.328
97153.266 61158.555 26063.672 554142.313 50540.059
GPTQ 13051.469 1301.395 107.458 236596.891 3543.708
7214.328 2782.353 52.472 282245.188 1665.287
AWQ 199448.797 93036.517 81834.344 1044956.250 31391.543
229888.406 74735.836 88684.156 639158.313 32668.666
QuIP 65.593 11.232 6.536 70.518 26.632
11.972 8.417 5.869 27.326 9.586
Table 1 summarizes PPL results of various quantized models evaluated on WikiText2, comparing several bit-widths and different layer-wise PTQ methods, both with and without QEP. Additional C4 and PTB dataset results are provided in Appendix D.1, demonstrating consistent trends in the following. Our results indicate that incorporating QEP significantly enhances the performance of layer-wise PTQ, substantially reducing perplexity across nearly all tested methods and quantization levels. In medium-bit scenarios such as INT4 and INT3, where AWQ already exhibits strong performance, applying QEP yields further improvements. At 2-bit quantization, existing layer-wise PTQ methods based on linear quantization typically suffer severe PPL degradation, rendering deployment infeasible. However, QEP effectively mitigates this issue, making INT2 quantization achievable with practical perplexity levels. Notably, QEP-enhanced QuIP achieves state-of-the-art perplexity results among all tested layer-wise PTQ methods. Similar significant improvements are observed for RTN, GPTQ, and AWQ at INT2g32, INT2g64, and INT2g128 quantization levels; see Appendix D.1 for details.
Zero-shot tasks
Table 2: Zero-shot average accuracy ($\uparrow$) on ARC-Easy, PIQA, and StoryCloze for Llama models across three quantization settings.
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B Llama-3-8B Mistral-7B
FP16 - - 0.7601 0.7840 0.8014 0.7920 0.8056
INT4 RTN 0.6802 0.7160 0.7325 0.7643 0.7831
0.6844 0.7131 0.7343 0.7686 0.7921
GPTQ 0.6817 0.7134 0.7306 0.4812 0.7906
0.6795 0.7104 0.7308 0.7531 0.7904
AWQ 0.6832 0.7120 0.7257 0.7821 0.7956
0.6870 0.7126 0.7331 0.7879 0.7967
QuIP 0.6500 0.7248 0.7285 0.7872 0.7204
0.6920 0.7167 0.7311 0.7800 0.8012
INT3 RTN 0.4770 0.6082 0.6402 0.4560 0.6448
0.5802 0.6550 0.6939 0.5388 0.6963
GPTQ 0.6367 0.6747 0.7043 0.4891 0.7305
0.6549 0.6853 0.7078 0.5901 0.7422
AWQ 0.5840 0.6886 0.7209 0.7074 0.7534
0.6264 0.6916 0.7283 0.7216 0.7675
QuIP 0.6232 0.7034 0.7246 0.7433 0.7422
0.6804 0.7128 0.7273 0.7549 0.7933
INT2 RTN 0.4139 0.4283 0.4147 0.4183 0.4130
0.4199 0.4191 0.4145 0.4108 0.4084
GPTQ 0.4162 0.4222 0.4356 0.4116 0.4159
0.4263 0.4283 0.4714 0.4228 0.4148
AWQ 0.4213 0.4176 0.4129 0.4164 0.4177
0.4162 0.4165 0.4140 0.4150 0.4181
QuIP 0.4667 0.5945 0.6628 0.4600 0.5422
0.5926 0.6404 0.6998 0.5121 0.6858
We evaluate the zero-shot accuracy of quantized models on several tasks. Table 2 summarizes the average accuracy for the ArcE, PiQA, and SC datasets. Detailed results for each dataset are provided in Appendix D.2. Consistent with the perplexity results, QEP effectively improves existing layer-wise PTQ methods. Notably, the performance gains from QEP are especially pronounced with INT2 quantization. For Llama-2-70B, the QEP-enhanced QuIP at INT2 achieves performance comparable to RTN and GPTQ at INT3 quantization.
Table 3: Runtime comparison of the quantization process.
Runtime Llama-2
7B 13B 70B
GPTQ 14.9m 26.4m 2.9h
AWQ 13.6m 25.4m 2.4h
QEP + RTN 10.9m 19.6m 1.7h
Runtime
We examine the impact of computation time required for the correction term. Table 3 shows the processing time of each layer-wise PTQ. Since the quantization processing time for RTN is only a few seconds and thus negligible, the measured time for QEP+RTN is primarily due to computing the preprocessing of the correction term. This result indicates that calculating the QEP correction term requires significantly less computation time than other existing layer-wise PTQ quantization processes. Moreover, using the same calibration dataset for weight correction and quantization reduces preprocessing overhead by approximately one-half to one-third by reusing computational steps.
Robustness
Table 4: Perplexity relative to RTN on WikiText2, comparing GPTQ and QEP+RTN calibrated with C4, PTB, and WikiText2.
PPL to RTN ($\downarrow$) Calibration Dataset
C4 PTB WikiText2
GPTQ -0.25 +0.07 -0.46
QEP + RTN -0.33 -0.30 -0.49
As discussed in Section 5.3, our method adaptively controls propagation strength in Eq. (10) to mitigate overfitting to the calibration dataset. In this section, we empirically validate this approach. Table 4 compares the perplexity difference among QEP-enhanced RTN, GPTQ, and RTN when quantizing Llama-2-7B, evaluated on WikiText2 across various calibration datasets. Consistent with prior findings (lin2024awq), GPTQ exhibits significant sensitivity to the calibration dataset: it outperforms RTN on C4 and WikiText2 but experiences notable performance degradation on PTB. In contrast, QEP-enhanced RTN consistently improves performance across all calibration datasets, demonstrating robustness to distributional shifts. This highlights the effectiveness of propagation control in preventing overfitting to the calibration dataset.
7 Conclusion
We revisit the core design of layer-wise PTQ and identify a critical limitation: the exponential accumulation and growth of quantization errors across network layers. To address this issue, we propose QEP, a general framework that explicitly propagates and compensates for accumulated quantization errors. Extensive experiments demonstrate that QEP substantially improves performance, especially in low-bit quantization scenarios. These findings underscore that meaningful progress in layer-wise PTQ can still be made by revisiting fundamental strategies, complementing recent trends primarily centered around non-linear and block-wise quantization techniques. Integrating QEP with these advanced quantization methods in the future presents a promising approach toward achieving extreme compression, potentially exceeding QAT performance.
Limitations
QEP relies on a small calibration set, as in other layer-wise PTQ approaches, which makes performance sensitive to data quality; however, it overfits less than comparable methods such as GPTQ and AWQ. The method also introduces a per-layer propagation-strength parameter $\alpha_{l}$; Although this parameter is tunable, a fixed value, e.g., $\alpha_{l}=\nicefrac{{1}}{{2}}$, works well in most cases, and automatic learning of $\alpha_{l}$ is left for future work.
Appendix A Additional Related Work
Quantization error mitigation
Defensive Quantization (DQ) [lin2019defensive] mitigates error accumulation by introducing an orthogonality penalty to the weights, which reduces the correlation-driven amplification of quantization noise It also employs gradient-based quantization to optimize the positioning of quantization levels, further suppressing propagated error and enhancing robustness. In contrast, the QEP module serves as a plug-and-play component for general layer-wise quantization, explicitly modulating propagation strength under standard linear, gradient-free quantization methods, such as GPTQ, AWQ, and QuIP.
Relation to GPTAQ
GPTAQ [li2025gptaq] optimizes the same local objective but is closely related to GPTQ, and therefore does not easily generalize to other layer-wise PTQ methods such as AWQ and QuIP. In contrast, QEP adds the correction term in Eq. (6) directly to the pre-trained weights, enabling plug-and-play use with diverse PTQ algorithms and ensuring strong performance in low-bit regimes. Moreover, whereas GPTAQ offers no guarantee that its local optimization reduces global quantization error, our analysis provides such a guaranty. QEP also introduces a per-layer propagation-strength parameter, $\alpha_{l}$, which mitigates the overfitting often observed with GPTQ and GPTAQ.
Extreme low-bit layer-wise PTQ
In the ultra-low precision regime, the layer-wise PTQ pipeline described in the main text, such as GPTQ and AWQ with standard linear quantizers, often becomes inadequate. As a result, many methods adopt alternative formalisms that introduce additional degrees of freedom beyond naive rounding to avoid catastrophic quality degradation. Representative examples include SVID-based 1-bit parameterizations [xu2024onebit], token-adaptive mixtures of scaling factors [jo2024binarymos], and structured sparsity designed for extreme quantization [dong2025stbllm]. In parallel, binary-factor formats decompose each weight matrix into bit-packed sign factors with lightweight diagonal scaling, allowing inference to be largely driven by efficient 1-bit kernels [boza2025dbf, lee2025littlebit]. In these formats, optimization is more challenging: strictly binary variables and highly non-smooth objectives can make straight-through estimators brittle [long2021learning, ichikawa2025high, yin2019understanding], motivating the use of discrete optimization or robust relaxation-based solvers. Recent advances in controlled continuous relaxations and annealing-style objectives have emphasized promising techniques such as QQA and iSCO [ichikawa2025continuous, ichikawa2025optimization, sun2023revisiting, ichikawa2024controlling]. A natural direction for future work is to examine how these extreme low-bit approaches interact with our error-propagation perspective. For instance, we should investigate whether combining QEP with binary-factor or sparsity-based methods can further suppress cross-layer error growth while maintaining the strong INT2 performance observed here, yielding larger gains as precision approaches 1 bit.
Appendix B Additional Theoretical Results
This section presents proofs of the propositions stated in the main text along with additional theoretical analyzes. To avoid ambiguity, we will fix the following notation throughout this section:
Symbol Description
${\bm{X}}\in{\mathbb{R}}^{d_{1}\times m}$ Calibration dataset (input activations at layer 1)
${\bm{W}}_{l}\in{\mathbb{R}}^{n_{l}\times d_{l}}$ Full-precision weight matrix at layer $l$
$\widehat{{\bm{W}}}_{l}\in{\mathbb{Q}}^{n_{l}\times d_{l}}$ Quantized weight matrix at layer $l$
$\sigma_{l}(\cdot)$ Activation function at layer $l$
${\bm{X}}_{l+1}\coloneqq\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})$ Full-precision activations at layer $(l+1)$
$\widehat{{\bm{X}}}_{l+1}\coloneqq\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l})$ Quantized activations at layer $(l+1)$
${\bm{\delta}}_{l}\coloneqq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$ Quantization error matrix at layer $l$
${\bm{H}}_{l}\coloneqq{\bm{X}}_{l}{\bm{X}}_{l}^{\top}$ Empirical Hessian of a full-precision model at layer $l$
$\widehat{{\bm{H}}}_{l}\coloneqq\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$ Empirical Hessian of a quantized model at layer $l$
In the following section, we assume that ${\bm{H}}_{l}$ and $\widehat{{\bm{W}}}_{l}$ are invertible. This assumption is standard in existing layer-wise PTQ methods, which also use these inverse matrices. To ensure numerical stability, a diagonal matrix $\rho{\bm{I}}$, $\rho>0$ is commonly added to the Hessian when its inversion becomes numerically unstable. The subsequent analysis remains consistent and valid even when applying this stabilization procedure, which simply involves adding $\rho{\bm{I}}$ to the Hessian in the following derivations.
Throughout this section, we examine the first-order linear term in the weight perturbations $\{{\bm{E}}_{l}\}$. Concretely, we define each quantity as ${\bm{A}}={\bm{A}}^{(0)}+{\bm{A}}^{(1)}+{\bm{R}}$, where ${\bm{A}}^{(1)}$ collects all terms linear in $\{{\bm{E}}_{l}\}$ and the remainder satisfies $\|{\bm{R}}\|={\mathcal{O}}(\max_{k}\|{\bm{E}}_{k}\|_{2}^{2})$ as $\max_{k}\|{\bm{E}}_{k}\|_{2}\to 0$. This matches practice: with INT8 rounding $\|{\bm{E}}_{l}\|_{F}/\|{\bm{W}}_{l}\|_{F}=10^{-2}\sim 10^{-1}$, so quadratic terms are one order of magnitude smaller than any first–order contribution. Furthermore, The baseline PTQ and the QEP pipeline use the same quantiser configuration, hence they induce errors of the same order. We therefore write the same symbol ${\bm{E}}_{l}$ for the error matrix in either scheme; any difference is at most a few percent and does not affect first–order bounds.
B.1 Derivation of Proposition 5.1
This section presents detailed proofs of Proposition 5.1 stated in the main text.
Proof.
First, we rewrite the residual inside the Frobenius norm by using the following relationship: ${\bm{W}}_{l}{\bm{X}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+{\bm{W}}_{l}{\bm{\delta}}_{l}$. Thus, the objective can be expressed as follows:
$\left\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\right\|_{F}^{2}=\left\|({\bm{W}}_{l}-\widehat{{\bm{W}}}_{l})\widehat{{\bm{X}}}_{l}+{\bm{W}}_{l}{\bm{\delta}}_{l}\right\|_{F}^{2}.$ (12)
Since the objective is a strictly convex quadratic function of $\widehat{{\bm{W}}}_{l}$ when $\widehat{{\bm{H}}}_{l}$ is invertible, the stationary point is the unique minimizer. To find the minimizer $\widehat{{\bm{W}}}_{l}$, we set the gradient of the expression with respect to $\widehat{{\bm{W}}}_{l}$, equal to zero. Using standard matrix calculus, we find that the calculus for a stationary point is
$({\bm{W}}_{l}-\widehat{{\bm{W}}}_{l})\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}+{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}={\bm{0}}.$ (13)
By defining $\widehat{{\bm{H}}}_{l}\coloneqq\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$, the above condition can be rewritten as
$({\bm{W}}_{l}-\widehat{{\bm{W}}}_{l})\widehat{{\bm{H}}}_{l}=-{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}.$ (14)
Assuming $\widehat{{\bm{H}}}_{l}$ is invertible, we multiply both sides on the right by $\widehat{{\bm{H}}}_{l}^{-1}$, obtaining
${\bm{W}}_{l}-\widehat{{\bm{W}}}_{l}=-{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1},$ (15)
and hence
$\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}+{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1}.$ (16)
This closed-form expression is indeed the unique minimizer of the Frobenius norm objective, thus completing the proof. ∎
B.2 Quantization Error Accumulation
This section demonstrates that, under standard layer-wise PTQ, where each layer is quantized independently without considering downstream effects, the activation difference at the output layer, defined as ${\bm{\delta}}_{L}\coloneqq{\bm{X}}_{L}-\widehat{{\bm{X}}}_{L}$, grows exponentially with depth, to first order in the quantization noise, under mild conditions.
Proposition B.1.
For each layer $l=1,\dots,L+1$, the activation error can be expressed as follows:
${\bm{\delta}}_{l}=-\sum_{k=1}^{l-1}\left(\prod_{s=k+1}^{l-1}{\bm{J}}_{s}{\bm{W}}_{s}\right){\bm{J}}_{k}{\bm{E}}_{k}{\bm{X}}_{k}+\mathcal{O}\left(\max_{k\leq l-1}\|{\bm{E}}_{k}\|_{F}^{2}\right),$ (17)
where the empty product $\prod_{s=l}^{l-1}$ is defined to be the identity matrix, and ${\bm{E}}_{k}\coloneqq\widehat{{\bm{W}}}_{k}-{\bm{W}}_{k}$ represents the weight quantization error at layer $k$.
Proof.
Consider explicitly the activations at layer $l$ in both full-precision and quantized forms:
${\bm{X}}_{l}=\sigma_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1}),~~\widehat{{\bm{X}}}_{l}=\sigma_{l-1}(\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1}).$ (18)
By recursively applying this relation back to the first layer, we derive the activation difference ${\bm{\delta}}_{l}$ as
${\bm{\delta}}_{l}$ $={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$ (19)
$=\sigma_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1})-\sigma_{l-1}(\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1})$ (20)
$={\bm{J}}_{l-1}({\bm{W}}_{l-1}{\bm{X}}_{l-1}-\widehat{{\bm{W}}}_{l-1}\widehat{{\bm{X}}}_{l-1})+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2}\}\right)$ (21)
$={\bm{J}}_{l-1}\left[-{\bm{E}}_{l-1}{\bm{X}}_{l-1}+{\bm{W}}_{l-1}{\bm{\delta}}_{l-1}\right]+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2}\}\right)$ (22)
$=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}+{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{\delta}}_{l-1}+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{\delta}}_{l-1}^{2}\}\right).$ (23)
By explicitly expanding ${\bm{\delta}}_{l-1}$, we obtain
${\bm{\delta}}_{l-1}=-{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}+{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}+\mathcal{O}\left(\max\{{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}\}\right).$ (24)
Substituting this expression into the previous equation yields
${\bm{\delta}}_{l}$ $=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}$ (25)
$~~+{\bm{J}}_{l-1}{\bm{W}}_{l-1}\left[-{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}+{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}\right]+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}\}\right)$ (26)
$=-{\bm{J}}_{l-1}{\bm{E}}_{l-1}{\bm{X}}_{l-1}-{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{J}}_{l-2}{\bm{E}}_{l-2}{\bm{X}}_{l-2}$ (27)
$~~+{\bm{J}}_{l-1}{\bm{W}}_{l-1}{\bm{J}}_{l-2}{\bm{W}}_{l-2}{\bm{\delta}}_{l-2}+\mathcal{O}\left(\max\{{\bm{E}}_{l-1}^{2},{\bm{E}}_{l-2}^{2},{\bm{\delta}}_{l-2}^{2}\}\right).$ (28)
By recursively repeating this explicit expansion down to the first layer, we obtain the fully expanded form as follows, noting ${\bm{\delta}}_{1}={\bm{0}}$:
${\bm{\delta}}_{l}=-\sum_{k=1}^{l-1}\left(\prod_{s=k+1}^{l-1}{\bm{J}}_{s}{\bm{W}}_{s}\right){\bm{J}}_{k}{\bm{E}}_{k}{\bm{X}}_{k}+\mathcal{O}\left(\max_{k\leq l-1}\|{\bm{E}}_{k}\|^{2}\right),$ (29)
where the empty product for $s=l,\dots,l-1$ is defined as the identity matrix. ∎
Proposition B.2.
Assume each activation $\sigma_{l}:{\mathbb{R}}^{n_{l}\times m}\to{\mathbb{R}}^{n_{l}\times m}$ is $\gamma_{l}$-Lipschitz with respect to the Frobenius norm and satisfies $\sigma_{l}({\bm{0}})={\bm{0}}$:
$\|\sigma_{l}({\bm{U}})-\sigma_{l}({\bm{V}})\|_{F}\leq\gamma_{l}\|{\bm{U}}-{\bm{V}}\|_{F},~~\gamma_{l}>0.$ (30)
Let ${\bm{X}}_{1}=\widehat{{\bm{X}}}_{1}={\bm{X}}$; for $l=1,\dots,L-1$ define
${\bm{X}}_{l+1}=\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l}),~~\widehat{{\bm{X}}}_{l+1}=\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}),~~\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}+{\bm{E}}_{l},$ (31)
and ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$. Assume $\|{\bm{W}}_{l}\|_{2}>0$ for all $l=1,\dots,L-1$ and set
$G_{L-1}\coloneq\prod_{l=1}^{L-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2},~~r\coloneq\max_{1\leq k\leq L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}.$ (32)
Then the final activation mismatch satisfies the explicit bound
$\|{\bm{\delta}}_{L}\|_{F}\leq\bigl((1+r)^{L-1}-1\bigr)G_{L-1}\|{\bm{X}}\|_{F}.$ (33)
Proof.
Since $\sigma_{l}({\bm{0}})={\bm{0}}$ and $\sigma_{l}$ are $\gamma_{l}$-Lipschitz,
$\|{\bm{X}}_{l+1}\|_{F}=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}({\bm{0}})\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}.$ (34)
By induction,
$\|{\bm{X}}_{l}\|_{F}\leq\left(\prod_{t=1}^{l-1}\gamma_{t}\|{\bm{W}}_{t}\|_{2}\right)\|{\bm{X}}\|_{F}=G_{l-1}\|{\bm{X}}\|_{F},~~l\geq 1.$ (35)
Using Lipschitz continuity again,
$\|{\bm{\delta}}_{l+1}\|_{F}$ $=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l})\|_{F}$ (36)
$\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}.$ (37)
Since $\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}+{\bm{E}}_{l}$ and ${\bm{\delta}}_{l}={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$,
${\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}{\bm{X}}_{l}-({\bm{W}}_{l}+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}=-{\bm{E}}_{l}{\bm{X}}_{l}+\widehat{{\bm{W}}}_{l}{\bm{\delta}}_{l},$ (38)
hence, by the triangle inequality and $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$,
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\gamma_{l}\Bigl(\|{\bm{E}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}+\|\widehat{{\bm{W}}}_{l}\|_{2}\|{\bm{\delta}}_{l}\|_{F}\Bigr).$ (39)
By definition of $r$, $\|{\bm{E}}_{l}\|_{2}\leq r\|{\bm{W}}_{l}\|_{2}$. Additionally, $\|\widehat{{\bm{W}}}_{l}\|_{2}\leq\|{\bm{W}}_{l}\|_{2}+\|{\bm{E}}_{l}\|_{2}\leq(1+r)\|{\bm{W}}_{l}\|_{2}$. Combining these with Eq. (35) in Eq. (39) yields
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}\|_{2}\Bigl(rG_{l-1}\|{\bm{X}}\|_{F}+(1+r)\|{\bm{\delta}}_{l}\|_{F}\Bigr).$ (40)
Define the normalized quantity
$a_{l}\coloneq\frac{\|{\bm{\delta}}_{l}\|_{F}}{G_{l-1}\|{\bm{X}}\|_{F}},~~l\geq 1.$ (41)
Note $a_{1}=0$ because ${\bm{\delta}}_{1}={\bm{0}}$. Dividing the previous inequality by $G_{l}\|{\bm{X}}\|_{F}$ (where $G_{l}=G_{l-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2}$) yields
$a_{l+1}\leq r+(1+r)a_{l}.$ (42)
Let $b_{l}\coloneq a_{l}+1$. Then $b_{l+1}\leq(1+r)b_{l}$ and $b_{1}=1$; hence $b_{l}\leq(1+r)^{l-1}$; therefore
$a_{l}\leq(1+r)^{l-1}-1.$ (43)
Taking $l=L$ gives Eq. (33). ∎
Proposition B.3.
Assume each activation $\sigma_{l}:{\mathbb{R}}^{n_{l}\times m}\to{\mathbb{R}}^{n_{l}\times m}$ is $\gamma_{l}$-Lipschitz with respect to the Frobenius norm and satisfies $\sigma_{l}({\bm{0}})={\bm{0}}$:
$\|\sigma_{l}({\bm{U}})-\sigma_{l}({\bm{V}})\|_{F}\leq\gamma_{l}\|{\bm{U}}-{\bm{V}}\|_{F}.$ (44)
Assume moreover that $\sigma_{l}$ is Fréchet differentiable at the full-precision pre-activation ${\bm{Y}}_{l}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}$ for each $l=1,\dots,L-1$. Let ${\bm{X}}_{1}=\widehat{{\bm{X}}}_{1}={\bm{X}}$ and for $l=1,\dots,L-1$ define
${\bm{X}}_{l+1}=\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l}),~~\widehat{{\bm{X}}}_{l+1}=\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}),~~\widehat{{\bm{W}}}_{l}={\bm{W}}_{l}+{\bm{E}}_{l}.$ (45)
Assume $\|{\bm{W}}_{k}\|_{2}>0$ for $k=1,\dots,L-1$ and define
${\bm{G}}_{L-1}\coloneq\prod_{l=1}^{L-1}\gamma_{l}\|{\bm{W}}_{l}\|_{2},~~r\coloneq\max_{1\leq k\leq L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}.$ (46)
Define the first-order component ${\bm{\delta}}_{L}^{(1)}$ as follows. For $t\in{\mathbb{R}}$ let $\widehat{{\bm{W}}}_{l}(t)\coloneq{\bm{W}}_{l}+t{\bm{E}}_{l}$ and define $\widehat{{\bm{X}}}_{1}(t)\coloneq{\bm{X}}$, $\widehat{{\bm{X}}}_{l+1}(t)\coloneq\sigma_{l}(\widehat{{\bm{W}}}_{l}(t)\widehat{{\bm{X}}}_{l}(t))$. Let ${\bm{\delta}}_{l}(t)\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}(t)$. Then ${\bm{\delta}}_{L}^{(1)}$ is defined by the derivative
${\bm{\delta}}_{L}^{(1)}\coloneq\left.\frac{d}{dt}\right|_{t=0}{\bm{\delta}}_{L}(t).$ (47)
Under these assumptions,
$\|{\bm{\delta}}_{L}^{(1)}\|_{F}\leq(L-1)r{\bm{G}}_{L-1}\|{\bm{X}}\|_{F}.$ (48)
In particular, if $\gamma_{l}\|{\bm{W}}_{l}\|_{2}\leq 1+\varepsilon$ for all $l$, then
$\|{\bm{\delta}}_{L}^{(1)}\|_{F}\leq(L-1)r(1+\varepsilon)^{L-1}\|{\bm{X}}\|_{F}.$ (49)
Proof.
Since $\sigma_{l}({\bm{0}})={\bm{0}}$ and $\sigma_{l}$ is $\gamma_{l}$-Lipschitz,
$\|{\bm{X}}_{l+1}\|_{F}=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}({\bm{0}})\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}.$ (50)
By induction,
$\|{\bm{X}}_{l}\|_{F}\leq\left(\prod_{t=1}^{l-1}\gamma_{t}\|{\bm{W}}_{t}\|_{2}\right)\|{\bm{X}}\|_{F},~~l\geq 1.$ (51)
Fix $l\in\{1,\dots,L-1\}$. By assumption, $\sigma_{l}$ is Fréchet differentiable at ${\bm{Y}}_{l}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}$. Let ${\bm{J}}_{l}\coloneq D\sigma_{l}({\bm{Y}}_{l})$ denote its Fréchet derivative. Because $\sigma_{l}$ is $\gamma_{l}$-Lipschitz, the operator norm of ${\bm{J}}_{l}$, induced by the Frobenius norm, satisfies
$\|{\bm{J}}_{l}\|_{\mathrm{op}}\leq\gamma_{l}.$ (52)
Indeed, for any $\bm{H}$},
$\|{\bm{J}}_{l}[{\bm{H}}]\|_{F}=\lim_{t\to 0}\frac{\|\sigma_{l}({\bm{Y}}_{l}+t{\bm{H}})-\sigma_{l}({\bm{Y}}_{l})\|_{F}}{|t|}\leq\lim_{t\to 0}\frac{\gamma_{l}\|t{\bm{H}}\|_{F}}{|t|}=\gamma_{l}\|{\bm{H}}\|_{F}.$ (53)
Now consider $\widehat{{\bm{X}}}_{l+1}(t)=\sigma_{l}(\widehat{{\bm{W}}}_{l}(t)\widehat{{\bm{X}}}_{l}(t))$. Since matrix multiplication is smooth and $\sigma_{l}$ is differentiable at ${\bm{Y}}_{l}$, the chain rule gives
$\left.\frac{d}{dt}\right|_{t=0}\widehat{{\bm{X}}}_{l+1}(t)={\bm{J}}_{l}\!\left[{\bm{E}}_{l}{\bm{X}}_{l}+{\bm{W}}_{l}\left.\frac{d}{dt}\right|_{t=0}\widehat{{\bm{X}}}_{l}(t)\right].$ (54)
Because ${\bm{\delta}}_{l}(t)={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}(t)$, we have $\left.\frac{d}{dt}\right|_{t=0}\widehat{{\bm{X}}}_{l}(t)=-{\bm{\delta}}_{l}^{(1)}$. Therefore,
${\bm{\delta}}_{l+1}^{(1)}=-{\bm{J}}_{l}[{\bm{E}}_{l}{\bm{X}}_{l}-{\bm{W}}_{l}{\bm{\delta}}_{l}^{(1)}]=-{\bm{J}}_{l}({\bm{E}}_{l}{\bm{X}}_{l})+{\bm{J}}_{l}({\bm{W}}_{l}{\bm{\delta}}_{l}^{(1)}),~~{\bm{\delta}}_{1}^{(1)}={\bm{0}}.$ (55)
Taking Frobenius norms, using Eq. (52) and $\|{\bm{A}}{\bm{B}}\|_{F}\leq\|{\bm{A}}\|_{2}\|{\bm{B}}\|_{F}$,
$\|{\bm{\delta}}_{l+1}^{(1)}\|_{F}\leq\gamma_{l}\|{\bm{E}}_{l}\|_{2}\|{\bm{X}}_{l}\|_{F}+\gamma_{l}\|{\bm{W}}_{l}\|_{2}\|{\bm{\delta}}_{l}^{(1)}\|_{F}.$ (56)
Define $a_{l}\coloneq\|{\bm{\delta}}_{l}^{(1)}\|_{F}$. From Eq. (56) and $a_{1}=0$, straightforward induction yields
$a_{L}\leq\sum_{k=1}^{L-1}\left(\prod_{s=k+1}^{L-1}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\gamma_{k}\|{\bm{E}}_{k}\|_{2}\|{\bm{X}}_{k}\|_{F}.$ (57)
Apply Eq. (51) to $\|{\bm{X}}_{k}\|_{F}$ and factor out ${\bm{G}}_{L-1}$:
$\|{\bm{\delta}}_{L}^{(1)}\|_{F}\leq{\bm{G}}_{L-1}\left(\sum_{k=1}^{L-1}\frac{\|{\bm{E}}_{k}\|_{2}}{\|{\bm{W}}_{k}\|_{2}}\right)\|{\bm{X}}\|_{F}\leq(L-1)r{\bm{G}}_{L-1}\|{\bm{X}}\|_{F},$ (58)
which is Eq. (48). If additionally $\gamma_{l}\|{\bm{W}}_{l}\|_{2}\leq 1+\varepsilon$ for all $l$, then ${\bm{G}}_{L-1}\leq(1+\varepsilon)^{L-1}$, proving Eq. (49). ∎
Proposition B.4.
Consider the 1-dimensional network ($d_{l}=n_{l}=1$) with $\sigma_{l}(z)=z$. Let $W_{l}=1+\varepsilon$ for all $l=1,\dots,L-1$ with $\varepsilon>0$, and let the input be $X=C>0$. Choose quantized weights
$\widehat{W}_{l}=W_{l}+E_{l},~~E_{l}\equiv c_{E}>0,~~l=1,\dots,L-1.$ (59)
Then, for all $L\geq 2$, the exact activation mismatch at layer $L$ satisfies
$|\delta_{L}|=|X_{L}-\widehat{X}_{L}|\geq(L-1)c_{E}C(1+\varepsilon)^{L-2}.$ (60)
In particular,
$|\delta_{L}|\geq\frac{c_{E}C}{1+\varepsilon}(1+\varepsilon)^{L-1}.$ (61)
Proof.
Since $\sigma_{l}$ is the identity map, we have
$X_{L}=(1+\varepsilon)^{L-1}C,~~\widehat{X}_{L}=(1+\varepsilon+c_{E})^{L-1}C.$ (62)
Hence
$|\delta_{L}|=C\left|(1+\varepsilon+c_{E})^{L-1}-(1+\varepsilon)^{L-1}\right|.$ (63)
Apply the mean value theorem to $f(t)=t^{L-1}$ on the interval $1+\varepsilon,\;1+\varepsilon+c_{E}]$: there exists $\xi\in(1+\varepsilon,1+\varepsilon+c_{E})$ such that
$(1+\varepsilon+c_{E})^{L-1}-(1+\varepsilon)^{L-1}=f^{\prime}(\xi)c_{E}=(L-1)\xi^{L-2}c_{E}.$ (64)
Since $\xi\geq 1+\varepsilon$, we obtain
$|\delta_{L}|\geq(L-1)c_{E}(1+\varepsilon)^{L-2}C,$ (65)
which proves Eq. (60). Finally, Eq. (61) follows from $(1+\varepsilon)^{L-2}=\nicefrac{{1}}{{1+\varepsilon}}(1+\varepsilon)^{L-1}$ and $L-1\geq 1$ for $L\geq 2$. ∎
B.3 Derivation of Theorem 5.2 and Corollary 5.4
This section presents a rigorous statement and proof of Theorem 5.2 and Corollary 5.4. We first formally restate Theorem 5.2 below.
Theorem B.5.
Consider an $L$-layer network
${\bm{X}}_{1}={\bm{X}},~~{\bm{X}}_{l+1}=\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l}),~~l=1,\dots,L.$ (66)
Assume each $\sigma_{l}$ is $\gamma_{l}$-Lipschitz with respect to $\|\cdot\|_{F}$ and satisfies $\sigma_{l}({\bm{0}})={\bm{0}}$. Let the quantized forward pass be
$\widehat{{\bm{X}}}_{1}={\bm{X}},~~\widehat{{\bm{X}}}_{l+1}=\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}).$ (67)
Define the activation mismatch as ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$.
Fix any matrices ${\bm{E}}_{l}$ and set
$\widehat{{\bm{W}}}_{l}^{\mathrm{BASE}}\coloneq{\bm{W}}_{l}+{\bm{E}}_{l}.$ (68)
For QEP, define each $l$
${\bm{W}}_{l}^{\ast}(\alpha_{l})\coloneq{\bm{W}}_{l}+\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1},~~\alpha_{l}\in[0,1],$ (69)
and set
$\widehat{{\bm{W}}}_{l}^{\mathrm{QEP}}\coloneq{\bm{W}}_{l}^{\ast}(\alpha_{l})+{\bm{E}}_{l}.$ (70)
Define the per-layer pre-activation residuals
${\bm{R}}_{l}^{M}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}^{M}\widehat{{\bm{X}}}_{l}^{M},~~M\in\{\mathrm{BASE},\mathrm{QEP}\},$ (71)
the global Lipschitz upper bound on the output mismatch
$\mathcal{U}^{M}\coloneq\sum_{k=1}^{L}\left(\prod_{s=k+1}^{L}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\gamma_{k}\|{\bm{R}}_{k}^{M}\|_{F}.$ (72)
Then for every choice of $\{\alpha_{l}\}_{l=1}^{L}\subset[0,1]$,
$\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}}.$ (73)
Consequently,
$\|{\bm{\delta}}_{L+1}^{\mathrm{QEP}}\|_{F}\leq\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}},~~\|{\bm{\delta}}_{L+1}^{\mathrm{BASE}}\|_{F}\leq\mathcal{U}^{\mathrm{BASE}}.$ (74)
Proof.
Fix a method $M\in\{\mathrm{BASE},\mathrm{QEP}\}$. By Lipschitz continuity,
$\|{\bm{\delta}}_{l+1}^{M}\|_{F}=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}(\widehat{{\bm{W}}}_{l}^{M}\widehat{{\bm{X}}}_{l}^{M})\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}^{M}\widehat{{\bm{X}}}_{l}^{M}\|_{F}=\gamma_{l}\|{\bm{R}}_{l}^{M}\|_{F}.$ (75)
Iterating this inequality through the remaining layers yields
$\|{\bm{\delta}}_{L+1}^{M}\|_{F}\leq\sum_{k=1}^{L}\left(\prod_{s=k+1}^{L}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\gamma_{k}\|{\bm{R}}_{k}^{M}\|_{F}=\mathcal{U}^{M}.$ (76)
It remains to show $\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}}$. We prove a stronger per-layer inequality:
$\|{\bm{R}}_{l}^{\mathrm{QEP}}\|_{F}\leq\|{\bm{R}}_{l}^{\mathrm{BASE}}\|_{F},~~\forall l.$ (77)
Fix $l$ and write $\widehat{{\bm{X}}}_{l}\coloneq\widehat{{\bm{X}}}_{l}^{\mathrm{QEP}}$ and ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$. Define the orthogonal projection
${\bm{P}}_{l}\coloneq\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l},$ (78)
which satisfies ${\bm{P}}_{l}^{2}={\bm{P}}_{l}$ and ${\bm{P}}_{l}^{\top}={\bm{P}}_{l}$.
By the construction of ${\bm{W}}_{l}^{\ast}(\alpha_{l})$, we have the exact identity
${\bm{W}}_{l}^{\ast}(\alpha_{l})\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}.$ (79)
Therefore, using $\widehat{{\bm{W}}}_{l}^{\mathrm{QEP}}={\bm{W}}_{l}^{\ast}(\alpha_{l})+{\bm{E}}_{l}$,
${\bm{R}}_{l}^{\mathrm{QEP}}$ $={\bm{W}}_{l}{\bm{X}}_{l}-({\bm{W}}_{l}^{\ast}(\alpha_{l})+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}$ (80)
$={\bm{W}}_{l}(\widehat{{\bm{X}}}_{l}+{\bm{\delta}}_{l})-{\bm{W}}_{l}^{\ast}(\alpha_{l})\widehat{{\bm{X}}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}$ (81)
$={\bm{W}}_{l}{\bm{\delta}}_{l}-\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}$ (82)
$={\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha_{l}{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}.$ (83)
Consider the BASE residual at the same layer evaluated on the same $\widehat{{\bm{X}}}_{l}$:
$\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}-({\bm{W}}_{l}+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}{\bm{\delta}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}.$ (84)
Hence
${\bm{R}}_{l}^{\mathrm{QEP}}=\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}-\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}.$ (85)
Since ${\bm{P}}_{l}$ is an orthogonal projection and $0\leq\alpha_{l}\leq 1$, Lemma B.8 implies
$\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha_{l}{\bm{P}}_{l})\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}\|_{F}.$ (86)
Therefore, by the triangle inequality,
$\|{\bm{R}}_{l}^{\mathrm{QEP}}\|_{F}=\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha_{l}{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}\leq\|\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}\|_{F}.$ (87)
Finally, since $\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}$ is the BASE activation produced by the BASE recursion, ${\bm{R}}_{l}^{\mathrm{BASE}}$ is exactly $\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}$ evaluated at $\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}$. Thus, taking $\widehat{{\bm{X}}}_{l}=\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}$ in the above inequality yields Eq. (77). Summing with nonnegative weights $(\prod_{s=k+1}^{L}\gamma_{s}\|{\bm{W}}_{s}\|_{2})\gamma_{k}$ yields $\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}}$, thus completing the proof. ∎
We further demonstrate that the final quantization error decreases monotonically as each propagation strength parameter $\alpha_{l}$ approaches $1$.
Corollary B.6.
Fix the layer index $l\in\{1,\dots,L\}$ and assume that $\widehat{{\bm{H}}}_{l}\coloneq\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$ is invertible. Let the activation mismatch be ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$. Define the orthogonal projection
${\bm{P}}_{l}\coloneq\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}\in{\mathbb{R}}^{m\times m}.$ (88)
For $\alpha\in[0,1]$, define the QEP corrected weight in the continuous domain as
${\bm{W}}_{l}^{\ast}(\alpha)\coloneq{\bm{W}}_{l}+\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1}.$ (89)
Let $\widehat{{\bm{W}}}_{l}(\alpha)\coloneq{\bm{W}}_{l}^{\ast}(\alpha)+{\bm{E}}_{l}$ be defined by a fixed matrix ${\bm{E}}_{l}$. Then, for any $0\leq\alpha^{\prime}\leq\alpha\leq 1$,
$\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha^{\prime}{\bm{P}}_{l})\|_{F}.$ (90)
Moreover, the pre-activation residual satisfies the exact identity
${\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}(\alpha)\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l},$ (91)
and hence the following upper bound is also monotone in $\alpha$:
$\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}(\alpha)\widehat{{\bm{X}}}_{l}\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})\|_{F}+\|{\bm{E}}_{l}\|_{2}\|\widehat{{\bm{X}}}_{l}\|_{F}.$ (92)
Proof.
First, ${\bm{P}}_{l}$ is an orthogonal projection. Indeed,
${\bm{P}}_{l}^{\top}=\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}={\bm{P}}_{l},$ (93)
and
${\bm{P}}_{l}^{2}=\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\underbrace{\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}}_{=\widehat{{\bm{H}}}_{l}}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}={\bm{P}}_{l}.$ (94)
Thus, Lemma B.8 applies with ${\bm{Z}}\coloneq{\bm{W}}_{l}{\bm{\delta}}_{l}$ and ${\bm{P}}\coloneq{\bm{P}}_{l}$, yielding
$\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha^{\prime}{\bm{P}}_{l})\|_{F},$ (95)
which proves Eq. (90).
Next, using the definitions of ${\bm{W}}_{l}^{\ast}(\alpha)$ and $\widehat{{\bm{H}}}_{l}^{-1}=(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}$,
${\bm{W}}_{l}^{\ast}(\alpha)\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}.$ (96)
Therefore, since $\widehat{{\bm{W}}}_{l}(\alpha)={\bm{W}}_{l}^{\ast}(\alpha)+{\bm{E}}_{l}$ and ${\bm{X}}_{l}=\widehat{{\bm{X}}}_{l}+{\bm{\delta}}_{l}$,
${\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}(\alpha)\widehat{{\bm{X}}}_{l}$ $={\bm{W}}_{l}(\widehat{{\bm{X}}}_{l}+{\bm{\delta}}_{l})-({\bm{W}}_{l}^{\ast}(\alpha)+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}$ (97)
$={\bm{W}}_{l}{\bm{\delta}}_{l}-\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}$ (98)
$={\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l},$ (99)
which is Eq. (91). Finally, Eq. (92) follows from the triangle inequality and $\|{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}\leq\|{\bm{E}}_{l}\|_{2}\|\widehat{{\bm{X}}}_{l}\|_{F}$. ∎
B.4 Relationship of QEP Correction and Ridge Regularization
We formally establish a rigorous mathematical connection between the Quantization Error Propagation (QEP) correction parameter $\alpha_{l}$ and the ridge regularization parameter $\lambda$. We show that tuning the QEP parameter $\alpha_{l}$ is equivalent to adjusting the strength of ridge regularization with parameter $\lambda$. We prove the monotone inverse relationship between these two parameters.
Proposition B.7.
The QEP update with mixing factor $\alpha_{l}\in[0,1]$ is
$\widehat{{\bm{W}}}_{l}^{\ast}(\alpha_{l})={\bm{W}}_{l}({\bm{I}}+\alpha{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1}).$ (100)
The unique minimizer of the ridge objective
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{R}}^{n_{l}\times d_{l}}}f(\widehat{{\bm{W}}}_{l}),~~f(\widehat{{\bm{W}}}_{l})=\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}^{2}+\lambda_{l}\|{\bm{W}}_{l}-\widehat{{\bm{W}}}_{l}\|_{F}^{2},~~\lambda_{l}\geq 0,$ (101)
equals
$\widehat{{\bm{W}}}^{\ast}_{l}(\lambda_{l})={\bm{W}}_{l}\left({\bm{I}}+{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}})^{-1}\right).$ (102)
Let the positive definite matrices be
${\bm{G}}(\alpha_{l})\coloneqq\alpha\hat{{\bm{H}}}_{l}^{-1},~~~{\bm{R}}(\lambda_{l})\coloneqq(\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}})^{-1}.$ (103)
Then
$\alpha_{1}\leq\alpha_{2}\Rightarrow{\bm{G}}(\alpha_{1})\preceq{\bm{G}}(\alpha_{2}),~~\lambda_{1}\leq\lambda_{2}\Rightarrow{\bm{R}}(\lambda_{1})\succeq{\bm{R}}(\lambda_{2}),$ (104)
and the scalar mapping as follows:
$\alpha(\lambda)\coloneqq\frac{1}{d}\mathrm{Tr}\widehat{{\bm{H}}}_{l}{\bm{R}}(\lambda)=\frac{1}{d_{l}}\sum_{i=1}^{d_{l}}\frac{\gamma_{i}}{\gamma_{i}+\lambda_{l}}$ (105)
with $\gamma_{1}\geq\cdots\geq\gamma_{d_{1}}>0$ the eigenvalues of $\widehat{{\bm{H}}}_{l}$, is strictly decreasing, satisfies $\alpha(0)=1$ and $\lim_{\lambda\to\infty}\alpha(\lambda)=0$, and obeys
$\mathrm{Tr}\widehat{{\bm{H}}}_{l}{\bm{G}}(\alpha(\lambda))=\mathrm{Tr}\widehat{{\bm{H}}}_{l}{\bm{R}}(\lambda).$ (106)
Thus, decreasing $\lambda$ from $\infty$ to 0 corresponds to increasing $\alpha_{l}$ from 0 to $1$.
Proof.
A standard differential identity $\partial\|{\bm{A}}\|_{F}^{2}=2{\bm{A}}$ gives
$\nabla_{\widehat{{\bm{W}}}_{l}}f(\widehat{{\bm{W}}}_{l})=2\left(\widehat{{\bm{W}}}_{l}\widehat{{\bm{H}}}_{l}-{\bm{W}}_{l}{\bm{X}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\right)+2\lambda(\widehat{{\bm{W}}}_{l}-{\bm{W}}_{l}).$ (107)
Setting this gradient $\bm{0}$} yields
$\widehat{{\bm{W}}}_{l}(\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}})={\bm{W}}_{l}({\bm{X}}_{l}\widehat{{\bm{X}}}_{l}^{\top}+\lambda{\bm{I}}),$ (108)
and right multiplication by inverse of $\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}}$ produces Eq. (102). Convexity of $f$ ensures uniqueness.
Diagonalise $\widehat{{\bm{H}}}_{l}={\bm{U}}{\bm{\Gamma}}{\bm{U}}^{\top}$ with $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{d_{l}})$. Then ${\bm{G}}(\alpha_{l})={\bm{U}}\alpha_{l}{\bm{\Gamma}}^{-1}{\bm{U}}^{\top}$ has eigenvalues $\nicefrac{{\alpha_{l}}}{{\gamma_{l}}}$, which increase strictly with $\alpha_{l}$, while ${\bm{R}}(\lambda_{l})={\bm{U}}({\bm{\Gamma}}+\lambda{\bm{I}})^{-1}{\bm{U}}^{\top}$ has eigenvalues $\nicefrac{{1}}{{\gamma_{i}+\lambda}}$, which decrease with $\lambda_{l}$, which means Loewner relations follow directly.
Furthermore, the following equation holds:
$\alpha(\lambda)=\frac{1}{d_{1}}\mathrm{Tr}(\widehat{{\bm{H}}}_{l}{\bm{R}}(\lambda))=\frac{1}{d_{l}}\sum_{i=1}^{d_{1}}\frac{\gamma_{i}}{\gamma_{i}+\lambda}$ (109)
Each summand has derivative
$\frac{\partial}{\partial\lambda}\frac{\gamma_{i}}{\gamma_{i}+\lambda}=-\frac{\gamma_{i}}{(\gamma_{i}+\lambda)^{2}}<0,$ (110)
which means $\alpha^{\prime}(\lambda)<0,\forall\lambda\geq 0$. Thus, $\alpha(\cdot)$ is strictly decreasing on $0,\infty)$. One has
$\lim_{\lambda\to 0}\frac{\gamma_{i}}{\gamma_{i}+\lambda}=1,~~~\lim_{\lambda\to\infty}\frac{\gamma_{i}}{\gamma_{i}+\lambda}=0.$ (111)
$\alpha(\cdot)$ is strictly decreasing from $1$ to 0 and smooth on $0,\infty)$. Because $\alpha$ is continous, strictly decreasing, it is a bijection from $0,+\infty)$ onto $(0,1]$. By construction
$\mathrm{Tr}\widehat{{\bm{H}}}_{l}{\bm{G}}(\alpha(\lambda))=\mathrm{Tr}\widehat{{\bm{H}}}_{l}{\bm{R}}(\lambda).$ (112)
Thus, decreasing $\lambda$ from $\infty$ to 0 corresponds to increasing $\alpha_{l}$ from 0 to $1$.
B.5 Technical Lemma
Lemma B.8.
Let ${\bm{Z}}\in{\mathbb{R}}^{m\times n}$ be arbitrary, and let $P\in{\mathbb{R}}^{n\times n}$ be an orthogonal projection, i.e., ${\bm{P}}^{2}={\bm{P}}$ and ${\bm{P}}^{\top}={\bm{P}}$. For every pair $0\leq\alpha^{\prime}\leq\alpha\leq 1$,
$\|{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\|_{F}\leq\|{\bm{Z}}(I-\alpha^{\prime}{\bm{P}})\|_{F}\leq\|{\bm{Z}}\|_{F}.$ (113)
Proof.
Write $f(\alpha)\coloneqq\|{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\|_{F}^{2}$. Because ${\bm{P}}^{\top}={\bm{P}}$ and ${\bm{P}}^{2}={\bm{P}}$,
$f(\alpha)=\mathrm{Tr}\left[({\bm{I}}-\alpha{\bm{P}}){\bm{Z}}^{\top}{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\right]=\|{\bm{Z}}\|_{F}^{2}-2\alpha(1-\alpha)\underbrace{\mathrm{Tr}({\bm{Z}}^{\top}{\bm{Z}}{\bm{P}})}_{t\geq 0}.$ (114)
Thus, $f^{\prime}(\alpha)=-(2-\alpha)t\leq 0$ on $[0,1]$ indicates that $f(\alpha)$ is non-increasing. Taking square roots yields the first inequality in Eq. (113). Setting $\alpha^{\prime}=0$ yields the second inequality: $\|{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\|_{F}\leq\|{\bm{Z}}\|_{F}$. ∎
Appendix C Additional Implementation Details
C.1 Damping for Hessian
A standard numerical issue in PTQ arises when the Hessian matrix $\widehat{{\bm{H}}}_{l}$ is ill-conditioned or singular, rendering its inversion unstable or undefined. Following GPTQ [frantar2022gptq], we resolve this issue by employing a damping strategy that adds a small scalar value $\lambda$ to the diagonal elements of $\widehat{{\bm{H}}}_{l}$ to ensure positive definiteness. In our implementation, we set $\lambda$ to the mean of the diagonal elements of $\widehat{{\bm{H}}}_{l}$, providing a straightforward yet effective method to stabilize the inversion process.
Figure 3: Results averaged over 5 random seeds comparing QuIP with and without QEP across different quantization levels. Each subplot shows results for INT4, INT3, and INT2 quantization, respectively, with the horizontal axis indicating model size (7B, 13B, 70B). The top row reports perplexity on WikiText-2 (lower is better), while the bottom row shows the average of normalized accuracy scores on ARC (easy), PIQA, and StoryCloze benchmarks (higher is better), representing generalization capability. Error bars represent the standard error of the mean (SEM). Models using QEP-QuIP consistently outperform or match the performance of baseline QuIP, especially under more aggressive quantization (INT3 and INT2).
Appendix D Additional Experiments
D.1 Additional Perplexity Results
Due to space constraints, the main text reports perplexity results solely for the WikiText-2 dataset. Here, we provide additional results for PTB (Table 6) and C4 (Table 7), along with extended results for WikiText-2 (Table 5). These supplementary results further validate that QEP consistently enhances PTQ performance, particularly in low-bit quantization scenarios.
Table 5: Perplexities ($\downarrow$) on WikiText-2 for Llama-2 (7B, 13B, 70B) under eight quantization settings.
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B
INT4g128 RTN 5.726 4.984 3.463
5.687 4.966 3.431
GPTQ 5.698 4.987 3.419
5.609 4.969 3.416
AWQ 5.599 4.987 3.408
5.580 4.969 3.404
INT4 RTN 6.116 5.206 3.672
6.017 5.165 3.621
GPTQ 6.083 5.167 3.594
5.933 5.127 3.576
AWQ 5.831 5.064 3.484
5.756 5.041 3.479
INT3g128 RTN 6.662 5.518 3.978
6.330 5.412 3.882
GPTQ 6.411 5.459 3.880
6.160 5.358 3.838
AWQ 6.247 5.315 3.740
6.108 5.295 3.724
INT3 RTN 539.866 10.688 7.530
17.309 7.458 5.648
GPTQ 10.881 6.632 4.860
7.898 6.245 4.102
AWQ 15.299 6.448 4.362
11.131 6.092 4.103
INT2g32 RTN 90.692 10.563 6.802
12.249 7.920 5.869
GPTQ 12.023 8.394 5.621
9.245 7.362 5.445
AWQ 15887.204 106933.227 63663.707
51.874 80654.797 37096.516
INT2g64 RTN 431.595 26.220 10.312
19.371 9.917 6.992
GPTQ 278.302 11.584 6.546
14.737 8.685 6.030
AWQ 217111.860 121737.148 71703.781
241136.594 126944.578 74227.539
INT2g128 RTN 4270.828 122.063 27.268
35.291 12.779 8.799
GPTQ 43.915 16.653 8.123
17.886 19.952 6.825
AWQ 222344.250 122795.898 72446.680
247751.203 126813.172 74192.570
INT2 RTN 17783.918 51152.832 26077.172
97153.266 61158.555 26063.672
GPTQ 13051.469 1301.395 107.458
7214.328 2782.353 52.472
AWQ 199448.797 93036.517 81834.344
229888.406 74735.836 88684.156
Table 6: Perplexities ($\downarrow$) on PTB for Llama-2 (7B, 13B, 70B) under eight quantization settings. “N/A” denotes numerical overflow (NaN).
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B
INT4g128 RTN 61.750 53.835 24.146
47.798 49.503 24.604
GPTQ N/A 51.133 24.101
N/A 50.072 24.243
AWQ 43.894 53.863 24.525
40.445 55.345 24.554
INT4 RTN 82.641 60.749 23.545
50.168 53.117 23.346
GPTQ N/A 53.561 24.720
124291.961 53.537 24.149
AWQ 60.261 56.152 25.542
46.937 57.445 24.411
INT3g128 RTN 55.467 64.638 23.586
48.576 54.866 24.776
GPTQ N/A 57.079 24.091
N/A 62.083 24.092
AWQ 64.932 57.273 24.668
52.356 61.479 26.309
INT3 RTN 37167.801 294.802 64.002
5514.820 113.856 34.212
GPTQ 44807.926 106.715 27.839
N/A 81.117 27.469
AWQ 130.308 121.698 26.887
81.606 93.260 25.592
INT2g32 RTN 20280.412 262.244 63.428
1685.683 96.913 36.677
GPTQ 18292.635 152.169 29.163
N/A 110.507 30.465
AWQ 47850.137 60977.195 48520.398
3741.642 47591.414 20185.246
INT2g64 RTN 9252.538 551.510 153.528
1096.720 158.306 42.991
GPTQ N/A 275.949 37.024
N/A 187.477 37.384
AWQ 202939.484 113584.867 79866.031
220728.234 117658.867 82598.511
INT2g128 RTN 9685.755 1213.282 767.896
4462.478 207.651 63.806
GPTQ 10694.694 395.689 56.685
N/A 325.407 45.569
AWQ 202164.484 113784.242 80543.727
222388.375 117059.742 82493.251
INT2 RTN 31824.279 42619.883 26063.672
10824.680 55286.305 26077.172
GPTQ N/A 3868.426 2438.034
N/A 3850.578 4050.844
AWQ 183984.766 87673.695 90442.352
198744.750 62160.063 91939.883
Table 7: Perplexities ($\downarrow$) on C4 for Llama-2 (7B, 13B, 70B) under eight quantization settings.
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B
INT4g128 RTN 7.584 6.869 5.826
7.513 6.839 5.786
GPTQ 7.522 6.860 5.778
7.421 6.828 5.770
AWQ 7.443 6.840 5.772
7.416 6.829 5.767
INT4 RTN 8.165 7.146 6.012
7.945 7.067 5.947
GPTQ 7.866 7.069 5.905
7.719 6.998 5.880
AWQ 7.721 6.962 5.842
7.634 6.932 5.828
INT3g128 RTN 8.977 7.582 6.266
8.510 7.402 6.150
GPTQ 8.502 7.463 6.105
8.185 7.316 6.072
AWQ 8.300 7.310 6.036
8.105 7.264 6.019
INT3 RTN 524.279 13.883 10.886
21.436 10.284 8.202
GPTQ 11.780 8.826 7.067
9.950 8.429 6.869
AWQ 17.418 9.049 6.631
13.934 8.257 6.353
INT2g32 RTN 225.440 13.879 9.720
16.148 10.561 8.459
GPTQ 14.365 10.719 7.932
11.839 9.685 7.717
AWQ 9028.133 76591.883 57596.215
51.811 49645.738 33026.816
INT2g64 RTN 553.766 30.445 15.155
22.089 12.762 9.850
GPTQ 20.860 13.394 8.981
14.084 11.039 8.508
AWQ 164477.422 95241.625 64913.477
181582.719 98917.820 67203.359
INT2g128 RTN 4811.772 131.665 47.878
34.022 15.398 12.081
GPTQ 33.370 18.008 10.535
18.184 12.704 9.433
AWQ 168465.266 95617.305 65646.594
187329.625 98457.031 67248.492
INT2 RTN 28258.385 52642.387 24912.074
108424.680 71050.250 29042.623
GPTQ 3048.671 299.684 56.719
276.638 629.527 30.874
AWQ 156266.797 81233.602 73251.945
177576.750 64098.504 75607.211
D.2 Detailed Accuracy Results for Individual Tasks
Due to space limitations, the main text reports only the average accuracy across three tasks. Here, we provide task-specific accuracies for PIQA (Table 8), StoryCloze (Table 9), and ARC-Easy (Table 10), further confirming that QEP consistently improves layer-wise PTQ.
Table 8: Accuracy ($\uparrow$) on PIQA for Llama-2 (7B, 13B, 70B) under eight quantization settings.
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B
INT4g128 RTN 0.773 0.792 0.804
0.773 0.790 0.806
GPTQ 0.770 0.789 0.807
0.771 0.792 0.806
AWQ 0.768 0.790 0.807
0.764 0.791 0.810
INT4 RTN 0.763 0.789 0.811
0.767 0.788 0.812
GPTQ 0.755 0.789 0.804
0.761 0.787 0.811
AWQ 0.760 0.789 0.807
0.763 0.784 0.814
INT3g128 RTN 0.757 0.770 0.793
0.761 0.779 0.806
GPTQ 0.758 0.778 0.806
0.764 0.782 0.807
AWQ 0.760 0.780 0.805
0.765 0.780 0.805
INT3 RTN 0.563 0.705 0.724
0.677 0.752 0.764
GPTQ 0.720 0.757 0.783
0.745 0.770 0.791
AWQ 0.647 0.760 0.787
0.725 0.770 0.801
INT2g32 RTN 0.588 0.696 0.760
0.693 0.735 0.771
GPTQ 0.690 0.732 0.772
0.714 0.748 0.776
AWQ 0.568 0.505 0.503
0.702 0.514 0.501
INT2g64 RTN 0.597 0.614 0.714
0.676 0.710 0.748
GPTQ 0.647 0.705 0.745
0.677 0.713 0.765
AWQ 0.502 0.506 0.502
0.702 0.506 0.504
INT2g128 RTN 0.511 0.566 0.635
0.652 0.678 0.721
GPTQ 0.581 0.639 0.715
0.659 0.683 0.747
AWQ 0.501 0.505 0.503
0.501 0.507 0.503
INT2 RTN 0.509 0.493 0.499
0.510 0.506 0.510
GPTQ 0.500 0.509 0.511
0.493 0.507 0.544
AWQ 0.507 0.504 0.502
0.505 0.504 0.504
Table 9: Accuracy ($\uparrow$) on StoryCloze for Llama-2 (7B, 13B, 70B) under eight quantization settings.
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B
INT4g128 RTN 0.765 0.785 0.791
0.770 0.788 0.794
GPTQ 0.768 0.784 0.793
0.771 0.789 0.798
AWQ 0.777 0.782 0.792
0.777 0.785 0.798
INT4 RTN 0.756 0.777 0.796
0.763 0.777 0.798
GPTQ 0.765 0.776 0.794
0.766 0.775 0.792
AWQ 0.760 0.774 0.789
0.766 0.777 0.794
INT3g128 RTN 0.749 0.766 0.790
0.756 0.773 0.789
GPTQ 0.763 0.776 0.793
0.759 0.770 0.796
AWQ 0.761 0.767 0.795
0.761 0.782 0.795
INT3 RTN 0.546 0.669 0.738
0.672 0.728 0.776
GPTQ 0.722 0.752 0.780
0.745 0.766 0.782
AWQ 0.689 0.767 0.787
0.702 0.764 0.782
INT2g32 RTN 0.645 0.668 0.745
0.704 0.721 0.776
GPTQ 0.758 0.715 0.724
0.763 0.748 0.766
AWQ 0.660 0.511 0.516
0.703 0.570 0.569
INT2g64 RTN 0.607 0.617 0.718
0.670 0.696 0.766
GPTQ 0.654 0.686 0.756
0.712 0.720 0.758
AWQ 0.476 0.479 0.476
0.474 0.479 0.475
INT2g128 RTN 0.509 0.577 0.647
0.651 0.677 0.741
GPTQ 0.588 0.634 0.724
0.649 0.690 0.753
AWQ 0.475 0.478 0.476
0.475 0.478 0.476
INT2 RTN 0.468 0.491 0.482
0.488 0.487 0.482
GPTQ 0.485 0.501 0.539
0.514 0.513 0.589
AWQ 0.489 0.478 0.475
0.482 0.476 0.477
Table 10: Accuracy ($\uparrow$) on ARC-Easy for Llama-2 (7B, 13B, 70B) under eight quantization settings.
Bits Method QEP Llama-2-7B Llama-2-13B Llama-2-70B
INT4g128 RTN 0.554 0.567 0.596
0.540 0.572 0.596
GPTQ 0.531 0.573 0.586
0.521 0.579 0.592
AWQ 0.537 0.577 0.585
0.526 0.580 0.592
INT4 RTN 0.521 0.582 0.590
0.524 0.574 0.593
GPTQ 0.525 0.575 0.594
0.512 0.570 0.589
AWQ 0.529 0.572 0.580
0.532 0.577 0.591
INT3g128 RTN 0.528 0.569 0.575
0.517 0.556 0.572
GPTQ 0.521 0.568 0.580
0.515 0.568 0.569
AWQ 0.534 0.561 0.597
0.527 0.561 0.592
INT3 RTN 0.322 0.450 0.459
0.391 0.485 0.541
GPTQ 0.468 0.514 0.550
0.474 0.520 0.551
AWQ 0.416 0.539 0.588
0.452 0.540 0.602
INT2g32 RTN 0.339 0.445 0.533
0.426 0.474 0.557
GPTQ 0.421 0.481 0.506
0.441 0.486 0.547
AWQ 0.352 0.272 0.263
0.449 0.280 0.263
INT2g64 RTN 0.332 0.371 0.467
0.390 0.430 0.557
GPTQ 0.377 0.455 0.485
0.404 0.458 0.548
AWQ 0.266 0.270 0.262
0.265 0.270 0.263
INT2g128 RTN 0.269 0.253 0.395
0.376 0.407 0.479
GPTQ 0.338 0.383 0.443
0.367 0.418 0.508
AWQ 0.266 0.269 0.260
0.265 0.269 0.261
INT2 RTN 0.265 0.253 0.263
0.262 0.264 0.261
GPTQ 0.263 0.256 0.257
0.272 0.265 0.281
AWQ 0.267 0.270 0.262
0.262 0.270 0.261
D.3 Stability of QuIP Results Across Random Seeds
We assess the stability of QuIP-only experiments by averaging five independent runs per configuration. Model sizes, quantization levels, and benchmarks align with the main Experiments section. Figure 3 plots QuIP with or without QEP at three quantization levels. Each marker is the mean of five seeds, and the error bars show the standard error of the mean. The top row gives perplexity on WikiText 2; the bottom row reports mean normalized accuracy on ARC easy, PIQA, and StoryCloze. Seed-to-seed variation is small and does not change the main conclusions. QEP-QuIP keeps its advantage, especially at INT3 and INT2. The main text lists the best seed per configuration for consistency with past work. This appendix confirms that the gains are not seed-specific but robust and reproducible, supporting using QEP.
Table 11: WikiText-2 perplexity for LLaMA-2-7B at different bit-widths. NaN denotes divergence.
Method INT4 INT3 INT2
RTN+QEP 6.017 17.309 97153.266
GPTQ+QEP 5.933 7.898 7214.328
AWQ+QEP 5.756 11.131 229888.406
QuIP+QEP 5.753 6.154 11.972
OmniQuant 5.880 7.065 NaN
D.4 Comparison with OmniQuant Baseline
For completeness, we compare QEP-enhanced layer-wise PTQ with block-wise OmniQuant [shao2023omniquant] on LLaMA-2-7B using WikiText-2 perplexity; lower values indicate better performance. As shown in Table 11, QuIP+QEP achieves the lowest perplexity at INT4/INT3 and remains stable at INT2, while OmniQuant diverges. These findings align with recent PTQ benchmarks indicating OmniQuant’s underperformance relative to layer-wise PTQ [zhao2025benchmarking].
[…]